OpenStudy (anonymous):

Please Help! Solve using elimination( the addition method) {10/3x+4y=-4 {5x+6y=-6

OpenStudy (anonymous):

multiply the first one by 3 and the second one by -2 and then your x's cancel out and then combine ur 2 y's together

OpenStudy (anonymous):

okay. thanks! could you double check my answer once i get it?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay so when i went to mulitply the 10/3. do i mulitply the 10 and 3?

OpenStudy (anonymous):

because i got 10x+12y=-12 -10x-12y=12

OpenStudy (anonymous):

when you multiply the 10/3 by 3 the three's cancel each other out and you will be left with 10

OpenStudy (anonymous):

but since that gave you that 0=0 try eliminating the y's and see what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you sure that you wrote the problems down right because when you eliminate the y you still end up with the same equation both spots

OpenStudy (anonymous):

i still get 0=0

OpenStudy (anonymous):

yes i did.

OpenStudy (anonymous):

you should get x=0 and y=-1

OpenStudy (anonymous):

okay. can you explain how you got those numbers please?

OpenStudy (anonymous):

see the thing is with this equation you can have an unlimited amount of solutions so since you found that when deleting the x out of the equation you get 0 then if you sub the 0 back in for x, you should then get y=-1, but if you eliminate y out you get 0 then if you sub that 0 back in you will get an answer other than 0 for x

OpenStudy (anonymous):

oh okay. thank you for your help!

OpenStudy (anonymous):

yep no problem glad i could help you!